OpenStudy (anonymous):

the volume of an enclosed cylindrical can with radius r cm and height h cm is 128000 cm^3,show that th surface area of the can is A=2pir^2+256000/r.find the value for r to minimize the surface area(state the value of r in term of pi)

OpenStudy (anonymous):

give me a few minutes plz i m doing it now

OpenStudy (anonymous):

i dont know how i m going get it interms of pi but i got the answer with decimal places

OpenStudy (anonymous):

is r = 10185.9?

OpenStudy (anonymous):

the answer is cube root(256000/4pi

OpenStudy (anonymous):

ouch.. thats like 27.31

OpenStudy (anonymous):

yeah did u get the solution for it?

OpenStudy (anonymous):

umm nopez.. i got something quite big compared to that..

OpenStudy (anonymous):

what i did was i differentiate A =2pir^2 + 256000/r and let it equal to the volume

OpenStudy (anonymous):

after that i solve for r

OpenStudy (anonymous):

hurm... its ok :D maybe i should try other question first thanx for the help

OpenStudy (anonymous):

wait i think i know how do it

OpenStudy (anonymous):

give me 2 mins

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

got it~~~

OpenStudy (anonymous):

Umm i am gonna write it out properly with the equations give me sometime

OpenStudy (anonymous):

ok thanks :D

OpenStudy (anonymous):

\[\pi r^2h = 128000\] \[r = \sqrt{\frac{ 128000 }{ \pi h }}\] u differentiate the surface area and Substitute your r into the equation dA/dr and let it equals zero so it will be like this.. \[ 4\pi \sqrt{\frac{ 128000 }{ \pi h }}) - \frac{ 256000 }{(\sqrt{\frac{ 128000 }{ h }})^2 } =0\] The from here u solve for h and substitute back to ur first equation r = squareroot (128000//pi h) to solve for r

OpenStudy (anonymous):

about the r interms of pi.. u need a graphic calculator that can be in standard mode to do it

OpenStudy (anonymous):

alright thanks for the help :D

OpenStudy (anonymous):

no problemz :)